A plane, traveling against a 60 km/hr headwind, covers 2100km in the same time that it could travel 2400km with a 70km tailwind. What is the speed of the plane with no headwind or tailwind?
Answer by mananth(16949)

let speed of plane = xmph
time against = time with wind
time against = 2100/(x-60)
time with wind = 2400/(x+70)
2100/(x-60)= 2400/(x+70)
2100(x+70)=2400(x-60)
2100x+147000=2400x-144000
2400x-2100x=147000+144000
300x=294000
/300
x=960 mph